OpenStudy (kirbykirby):

Linear Algebra: If S = {(a, b) in R^2 | b>0} and we define addition to be (a, b) [+] (c, d) = (ad+bc, bd) I need to show that (a,b) [+] (c, d) belongs to a vector space. Is it sufficient to say that b>0 and d>0, so bd>0 ? Or do we have to say something about the ad+bc as well?

OpenStudy (amistre64):

therer are at least 3 conditions for a vector space

OpenStudy (amistre64):

closed under addition closed under multiplication contains a zero vector

OpenStudy (amistre64):

the rest of the conditions are consequences of these

OpenStudy (kirbykirby):

Oh yes I know what you mean. I meant I wa strying to show the first axiom. Saying that in genral, for vectors x', y' belonging to a vector space V, then x' + y' is in V

OpenStudy (amistre64):

i take it these are your generic vectors? <a,b> and <c,d>

OpenStudy (kirbykirby):

yes

OpenStudy (amistre64):

and the only conditions are R^2 meaning? the vectors themselves are in R^2? or that their component elements are in R^2 (which makes no sense if thats the case)

OpenStudy (kirbykirby):

I believe they mean the vectors themselves are in R^2

OpenStudy (amistre64):

then by the definition they gave of the additive rule would have to be parsed; your assumption the d>0 is superfluous to me

OpenStudy (kirbykirby):

well since (a, b) are in R^2, b>0, and if (c, d) is in R^2, then d>0 also no?

OpenStudy (amistre64):

no, d is any real value; only b is a positive number

OpenStudy (kirbykirby):

but aren't they saying that (a, b) is any vector in R^2?

OpenStudy (kirbykirby):

So I thought they meant the second component of any vector is >0

OpenStudy (amistre64):

they are saying that any vector in the space of the for <a,b> is in R^2; its a generic vector and all the forms of it are as well <a,b> + <c,d> ------------ <ad+bc , bd>

OpenStudy (amistre64):

since addition and multiplication are closed under the Reals the left and right side will be simply of the form <x,y> if we assume c.d to be the zero vector for addition <0,0> do we get <a,b> back?

OpenStudy (anonymous):

not to butt in, but do you know what this is?

OpenStudy (amistre64):

the zero vector is tends to be the easiest one to try to figure out

OpenStudy (amistre64):

this is openstudy :)

OpenStudy (anonymous):

lol

OpenStudy (amistre64):

i believe they re trying to define a "group"

OpenStudy (anonymous):

i meant this <a,b> + <c,d> ------------ <ad+bc , bd>

OpenStudy (kirbykirby):

but don't we have to restrict d>0, we can't have <0,0> right according to the restriction?

OpenStudy (kirbykirby):

o.O

OpenStudy (amistre64):

the binary operator is defined that way

OpenStudy (anonymous):

<0,0> is not the identity

OpenStudy (amistre64):

what it means? i dunno

OpenStudy (anonymous):

hint, \[\frac{a}{b}+\frac{c}{d}=...\]

OpenStudy (amistre64):

if c,d can be the inverse of a,b then we can define an identity from it perhaps?

OpenStudy (amistre64):

...thought I had an idea lol. im lost again

OpenStudy (anonymous):

adding fraction is what you are doing

OpenStudy (anonymous):

identity is <0,b>

OpenStudy (amistre64):

ahhh, ok

OpenStudy (anonymous):

well actually i guess you need so say something about equivalence classes, so maybe i should be quiet

OpenStudy (kirbykirby):

Wow this question seems more confusing that I even thought lol

OpenStudy (amistre64):

the closest ive come to learning eq classes is reading the word in a group theory book ....

OpenStudy (kirbykirby):

Oh I don't know much about group theory. We haven't done it yet. This is only my first linear algebra class =( lol

OpenStudy (amistre64):

spose the "vectors" look like this:\[\binom{a}{b}\] then the binary operation looks more like the addition of fractions

OpenStudy (anonymous):

typo there. the identity is <0,b> if you can do the following <0,b> + <c,d> = <bc,bd> now with fractions we can "cancel" the b and get that this is <c,d> making <0,b> the identity

OpenStudy (anonymous):

so what is missing from this set up is a definition of what <a,b> = <c,d> would be

OpenStudy (anonymous):

how would we know when two of these are equal with fractions we know it we know \[\frac{a}{b}=\frac{c}{d}\iff ad =bc\]

OpenStudy (kirbykirby):

We have do define was <a,b> = <c, d> is? All I have from the question is what the set S was :s

OpenStudy (kirbykirby):

to*

OpenStudy (amistre64):

:) sneaky little question indeed

OpenStudy (kirbykirby):

Are we allowed to say though that we are manipulating fractions, even though nothing is mentioned?

OpenStudy (anonymous):

then maybe i am reading too much in to it, maybe you should just check the vector space axioms. but if we don't know when two are equal, how would we know what the identity is?

OpenStudy (anonymous):

All we have to to is show that as long as b and d are positive integers then vectors (a,b) and (c,d) are in the subspace of S because it satisfies b+d>0

OpenStudy (anonymous):

Is that wrong?

OpenStudy (anonymous):

you need to know what the zero vector is, because it needs to be in your subspace

OpenStudy (anonymous):

and in order to know what the zero vector is, you need to know when two vectors are the same

OpenStudy (anonymous):

how do i know that 0 + v = v if i don't know what conditions give me v = v

OpenStudy (anonymous):

so the question is, what is the identity under this operation. with fractions it is easy, it is 0/b which translates to <0,b> in this case, but that is because we know that dc/bd = c/d

OpenStudy (kirbykirby):

hm ok I kinda see where you're getting at. But we can't say that once we add the zero vector on the right side, and obtain the left side, then naturally v = v? without needing to define it beforehand?

OpenStudy (amistre64):

@Romero how would showing that b+d>0 satisfy the question: I need to show that (a,b) [+] (c, d) belongs to: a vector space

OpenStudy (amistre64):

as long as we dont try to paint the problem with our biased math notions .. we should be good

OpenStudy (anonymous):

maybe i am reading too much into it the question says I need to show that (a,b) [+] (c, d) belongs to a vector space. and i am not really sure what this means "belongs to a vector space"

OpenStudy (amistre64):

d can = 0 so long as it doesnt mess up the program

OpenStudy (anonymous):

question is king of goofy does "a" belong to a vector space?

OpenStudy (amistre64):

lol, i was assuming the elements were composed of real numbers ...

OpenStudy (anonymous):

did the question come out of a book? is that the exact wording?

OpenStudy (anonymous):

sure but i mean how can you ask if something "is an element of a vector space" ??

OpenStudy (amistre64):

by satisfying the axioms that define a vector space

OpenStudy (anonymous):

but the question doesn't say "is this a vector space" which i could understand. it says I need to show that (a,b) [+] (c, d) BELONGS to a vector space.

OpenStudy (anonymous):

"I need to show that (a,b) [+] (c, d) belongs to a vector space." Do you have to do this. Is this what the problem is asking you or are you trying to see if it belongs to a vector space?

OpenStudy (kirbykirby):

Well they had a similar example in the book: they had S = {x in R | x > 0}. Define addition in S by x [+] y = xy. Prove that S is a vector space. 1) Show x [+] y is in S: x [+] y = xy. Since x >0 and y>0, then x[+]y is in S

OpenStudy (anonymous):

you have elements of R^2 that is ordered pairs of real numbers, and you have a definition of addition. so you could ask "does this set under this operation form a vector space?" that question would make sense

OpenStudy (amistre64):

right, which I read as; take the elements as defined in R^2 and perform a binary operation on them; is the results of that operation conform to the concept of a vector space; or is it only a subset and not a subspace?

OpenStudy (anonymous):

holy crap!

OpenStudy (anonymous):

i get what the problem is asking, show that it is closed!!

OpenStudy (anonymous):

It looks like you are making assumptions about the values of the variables. So can you answer the question with a specific example?

OpenStudy (amistre64):

specifics are never "proof"

OpenStudy (anonymous):

but unless i am senile that second question you asked is wrong, that does not form a vector space in a vector space a(x+ y) = ax + ay for all vectors x,y and scalers a

OpenStudy (anonymous):

Not trying to prove anything but rather show a contradiction.

OpenStudy (anonymous):

and it is certainly not true with "+" defined as multiplication, because no way is a(xy) = axay

OpenStudy (amistre64):

contradiction by example are always welcome ;)

OpenStudy (anonymous):

should be Well they had a similar example in the book: they had S = {x in R | x > 0}. Define addition in S by x [+] y = xy. Prove that S is a NOT vector space.

OpenStudy (kirbykirby):

Ohh wait. I'm so sorry. Later on the page, when they are proving the scalar multiplication stuff. They say "Define scalar multiplication to be c [*] x = x^c

OpenStudy (anonymous):

ok well i am sorry i butted in, i thought the question was "show this is a vector space" but what it looks like it is asking is "show that it is closed" so you have to show that you end up in the space, which you do since if b > 0 and d > 0 certainly bd > 0